# Advancing Identification method of Gamma-Ray Bursts with Data and Feature Enhancement

### v1

@@ 1 Introduction
Gamma-ray bursts are among the most energetic phenomena in the universe, releasing enormous amounts of energy in the form of gamma rays over remarkably short time frames. Despite decades of remarkable progress in observations and theoretical investigations, their origins remain elusive (Zhang2011CRP; KumarZhang2015PhR; Meszaros2019MmSAI; Peer2024Galax), making GRB identification and classification a critical focus of contemporary astrophysical research (lvhoujun2010ApJ; ZhangShuai2022ApJ; Sun2023Arxiv; Wang2024Arxiv). The rise of multi-messenger and multi-band astronomy further underscores the need for efficient GRB detection (Margutti2021ARAA; Rudolph2023ApJ): timely and accurate detection enables follow-up observations of afterglows, host galaxy localization, characterization of associated counterparts, and constraints on GRB intrinsic properties.
@@
GRB light curves often exhibit irregular, multi-peaked structures that encode rich physical information about their outbursts. While traditional classification criteria (e.g., duration, hardness) have been widely adopted, emerging evidence highlights the need for additional discriminants to address the ”iceberg effect”—a phenomenon where low-signal-to-noise ratio (low-SNR) GRBs are partially submerged in background noise, disproportionately impacting short-duration events (2014MNRAS.442.1922L; iceberg-effects_2022Moss). Traditional burst search algorithms identify GRBs by detecting signals in multi-time-bin, multi-energy-band light curves where the SNR exceeds a predefined threshold above the background (fermi_blind_search; fermi_target_search; search_grb_hxmt_cai). However, these methods face significant challenges, including accurate background estimation and optimal threshold selection, limiting their effectiveness in identifying faint or complex GRB events. These limitations motivate the development of advanced techniques to enhance GRB detection and classification accuracy.
@@
In recent years, machine learning (ML) has emerged as a experimental tool for GRB detection. For example, relate_search_GRB_DTW developed an ML algorithm for automated detection of GRB-like events using AstroSat/CZTI data, combining density-based spatial clustering with dynamic time warping to demonstrate ML’s versatility and robustness in GRB identification. Among ML architectures, convolutional neural networks (CNNs)—a specialized deep learning (DL) framework—excel in pattern recognition tasks (Alzubaidi2021ReviewOD; Taye2023TheoreticalUO), and have driven a paradigm shift in astrophysical data analysis by autonomously extracting features from multi-dimensional, multi-modal datasets. In GRB research, CNN-based approaches have achieved notable successes. DL_detect_GRB_intensity_map developed a real-time CNN pipeline for AGILE-GRID intensity maps, detecting 21 GRBs at 3 significance compared to only two detections via conventional SNR-based searches. DL_autodencoder_detect_grb used a CNN autoencoder to reconstruct background light curves and identify 72 previously uncataloged AGILE GRBs. Additionally, (ParmiggianiarX240402107P) and nn_search_long_faint_burst leveraged neural networks to predict background count rates and identify faint, long-duration GRBs.
@@
Notably, GRB identification relies on extracting meaningful features from light curves with complex temporal and spectral characteristics. While traditional 1D-CNNs effectively capture local temporal patterns, their inability to model spectral-temporal correlations limits their utility in GRB analysis—critical GRB features are jointly encoded in both domains (GRB-fft_2016Guidorzi; GRB-fft_2021Tarnopolski; GRB-fft_2024Zhou). To address this, recent time series classification advances have introduced architectures that explicitly model frequency information, including Wavelet-based decomposition (e.g., T-WaveNet (CNN_freq_liu_2020)), 2D temporal representations (e.g., TimesNet (CNN_freq_wu_2022)), hybrid Fourier-wavelet transformers (e.g., FEDformer (CNN_freq_zhou_2022), WFTNet (CNN_freq_liu_2024)), and adaptive spectral blocks (e.g., TSLANet (model_block_ASB)). These state-of-the-art approaches confirm that spectral-temporal modeling enhances feature extraction, underscoring the need for GRB identification methods that integrate both temporal dynamics and frequency-domain characteristics.
@@
Complementary to classification, dimensionality reduction techniques have provided valuable insights into GRB diversity. For instance, DownDimGRB_jesperen2020 used t-distributed stochastic neighbor embedding (t-SNE) to categorize Swift/BAT GRBs to separate supernova-associated and kilonova-associated GRBs into long and short categories. ML-downdim-GRB-EE_2023Garcia identified 7 new extended emission (EE) GRB candidates via dimensionality reduction. 2024MNRAS.532.1434Z and DownDimGRB_Negro2025 used Uniform Manifold Approximation and Projection (UMAP) or t-SNE to reveal GRB clusters that transcend traditional duration-based classification, including distinct subgroups of kilonova-related GRBs. Similarly, 2024MNRAS.527.4272C_Chen and DownDimGRB_Chen2025 demonstrated robust bimodal GRB clustering and refuted intermediate GRB classes using unsupervised ML and dimensionality reduction. Collectively, these studies highlight the utility of such techniques for analyzing burst similarities and refining classification schemes.
@@
Despite these advances, DL models face critical limitations in GRB research. First, data scarcity persists: the largest available GRB dataset (Fermi/GBM, 3900 events) remains insufficient for robust DL training (DL_identify_grb_by_peng), due to limited labeled observations, skewed distribution (underrepresentation of rare, short-duration, and low-SNR GRBs), and narrow coverage of physical properties (e.g., redshift, energy bands, progenitors). While data augmentation has been explored to mitigate this (DL_identify_grb_by_peng; DataAug_2019BigDataSurvey), existing datasets still lack the feature richness needed to avoid overfitting. Second, uncertainty quantification (UQ) is underdeveloped: model uncertainty stems from limited training data, the ”black-box” nature of CNNs, and observational noise (Nemani2023MSSP20510796N), yet UQ is critical for distinguishing high-confidence detections from ambiguous candidates—misclassification of which could lead to missed follow-up observations or erroneous inferences about progenitor origins.
@@ 2 Data augmentation and Dataset
The Fermi/GBM trigger system has been operational since 12 July 2008 (first detected GRB: GRB 080714B). For this study, we utilized Fermi/GBM data spanning from 14 July 2008 to 30 June 2024, including 3,905 original GRBs detected by NaI detectors. These events are manually verified and cataloged in the Fermi/GBM Burst Catalog111https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html. Adopting the GRB/non-GRB extraction protocol described in DL_identify_grb_by_peng, we updated the dataset by segmenting samples into 120-second time windows. To ensure complete coverage of the interval (with background included on both sides), 217 GRBs failing this criterion were excluded. For the remaining events, samples were generated based on the number of triggering NaI detectors, resulting in 6,189 primary GRB samples. Non-GRB samples—encompassing modulated background, electronic noise, and potential all-sky source signals—were systematically extracted from detectors’ quiescent (non-trigger) intervals, totaling 108,000. This balance between classes enhances model generalization to non-burst events.
Light curves for all samples were extracted at a 64 ms temporal resolution across 128 energy channels, then rebinned into 9 standard energy bands (25–50, 50–100, 50–300, 100–300, 100–500, 100–900, 300–500, 300–900, 500–900 keV). This segmentation aligns with typical GRB photon energy distributions and Fermi/GBM trigger criteria (fermi_gbm_fouth_catalog; search_grb_hxmt_cai), approximating the photon deposition ranges of NaI detectors (without energy response correction). Critically, it preserves energy-dependent features such as burst morphology evolution, light curve shape variations, and spectral lags—enriching the diversity of discriminative features for model training. To optimize input for machine learning, we applied per-band standardization to light curves (scaling features to zero mean and unit variance). This approach preserves spectral information via inter-band flux ratios, empirically outperforming normalization (scaling to [0,1]). While both methods unify feature scales to facilitate deep learning convergence, standardization is preferable for GRB characterization as it retains spectral integrity—essential for distinguishing GRBs from non-GRBs (e.g., via hardness evolution or spectral lags).
@@
Remove 4,648 synthetic samples with peak-SNR blew 2 , as excessive photon subtraction makes these resemble background samples.
@@
Dataset: We partition preprocessed samples into training, validation, and test sets chronologically to ensure temporal generalization. Dataset details are shown in Table 1. This chronological split alleviates temporal overfitting while preserving the data’s inherent flux distribution. We maintain 1:1 positive-to-negative sample ratios across all splits. Crucially, only the training set undergoes data augmentation, validation/test sets retain original observations (primary GRB samples and non-GRB samples) for unbiased performance evaluation.
@@ 3.1 Architecture of Neural Networks
Each sample in our dataset consists of light curves spanning 9 energy bands, which are naturally suited to be treated as time series data. This makes a 1D CNN the appropriate choice for feature extraction and classification. Our model adopts ResNet (model_resnet) as its backbone, which relies on a sequential stack of four convolutional units (Conv Units) to progressively extract hierarchical features from the input light curves. A key advantage of ResNet is its integration of residual connections, which effectively alleviate the gradient vanishing problem that commonly arises in deep neural networks and enable stable training of deeper architectures. To address the limitation that CNNs lack the ability to distinguish subtle frequency variations, we propose an Adaptive Frequency Feature Enhancement module, AFFE, which is integrated into the last part of each Conv Unit. This module explicitly enhances frequency-domain features to improve the model’s GRB identification capability, and it also adaptively weights and filters frequency components, enabling the model to focus on the most discriminative features for GRB recognition. This design is particularly critical for distinguishing GRBs from background noise, as frequency-domain analysis can uncover patterns that remain obscured in the time domain.
After the convolutional layers (including Conv Units with AFFE modules) complete feature extraction, the extracted high-dimensional features first undergo global average pooling (GAP). This step reduces the number of model parameters while preserving the representative information of features across the temporal dimension, thereby mitigating the risk of overfitting. A dropout layer with a 50% dropout rate is then inserted between the GAP layer and the subsequent fully connected (FC) layers. During training, this layer randomly deactivates half of the neurons to prevent the model from over-relying on specific features, further suppressing overfitting. Following the dropout layer, two FC layers map the pooled features to the final classification space and output the probability of the sample belonging to the GRB or non-GRB category. The overall architecture of our our ResNet-based model integrated with AFFE modules is shown in Figure 3, which includes the order of layers and the data flow.
@@ 3.2 Model training and analysis
Model parameters are initialized using the truncated normal distribution proposed by param_he_normal. A detailed analysis of hyperparameter choices-including batch size, learning rate, and number of convolutional layers-is conducted to evaluate their impact on the test set, following the approach in dl_technosignatures. Further details on hyperparameter selection, such as the number of convolutional kernels per unit, convolutional kernel sizes, and neuron counts in FC layers. The training configuration is as follows: batch size is set to 1024; cross-entropy loss is used to measure the discrepancy between predicted and true labels; and the Adam optimizer adam_optimiser is adopted to dynamically update model parameters and minimize loss during training. The initial learning rate is , with a scheduling mechanism that reduces the rate by half if validation loss does not decrease for 10 consecutive epochs. To prevent overfitting, early stopping terminates training if validation accuracy plateaus for 20 consecutive epochs. The patience value was selected after testing ranges of [10, 20, 40] epochs, and the model parameters from the epoch with the highest validation accuracy are retained as the final optimal parameters. The result of the optimizing choices are highlighted in bold in Table 2. The efficacy of selected hyperparameters is further validated by training curves, which demonstrate stable convergence and appropriate model capacity for the classification task.
@@
We conducted a comparative analysis involving the baseline ResNet architecture that commonly used for time series data, and ResNet variants integrated with existing feature enhancement modules that including the SE (Squeeze-and-Excitation block; model_block_SE), CBAM (Convolutional Block Attention Module; model_block_CBAM), ECA (Efficient Channel Attention module; model_block_ECA), and ASB (Adaptive Spectral Block; model_block_ASB modules), and our proposed ResNet-AFFE model. All models were trained and tested on our dataset using hyperparameters consistent with those reported in the original literature for each comparative module. This ensures a fair performance comparison, and the results of this comparison (for ResNet variants and our ResNet-AFFE model) on the test set are presented blew.
The models were implemented using the PyTorch framework, with training conducted on a single NVIDIA RTX-4090 GPU. Performance was evaluated using four standard classification metrics: Accuracy, Precision, Recall, and F1-score, consistent with the evaluation methodology reported in DL_identify_grb_by_peng. Accurately verifying these metrics and their associated prediction uncertainties is critical, as they quantify confidence in model predictions and provide a scientific basis for subsequent decision-making. During testing, we achieved robust uncertainty quantification for deep learning predictions via Monte Carlo Dropout (MCD, for references see model_uncertainty1; model_uncertainty2). In this procedure, all model parameters are kept fixed, and dropout layers between FC layers remain activated. For each forward pass, the dropout rate is randomly sampled from a uniform distribution over the interval [0.1, 0.5]. This stochastic forward pass is repeated 1000 times, which allows computation of the mean and standard deviation for each metric—thereby effectively estimating predictive uncertainty. The performance and uncertainty of all models are summarized in Table 3.
@@
Enhanced time domain feature extraction module: a model_block_SE, b model_block_CBAM, c model_block_ECA.
Enhanced frequency domain feature extraction module: d model_block_ASB.
@@
To examine whether ResNet-AFFE prioritizes burst-specific temporal signatures (e.g., intervals or energy-dependent flux peaks) over noise, we used Gradient-weighted Class Activation Mapping (Grad-CAM)—a mainstream visualization technique for addressing deep learning’s ”black box” problem. Grad-CAM generates gradient-weighted heatmaps to identify features most influential to model predictions (model_grad_cam), enhancing interpretability. For time-series light curves, these heatmaps pinpoint temporal segments driving classification outputs, replacing ambiguous ”feature importance” with intuitive visual evidence. We employ Grad-CAM to decode our model’s decision-making process for GRB classification. Comparative visualizations of discriminative temporal features between the baseline ResNet and ResNet-AFFE are presented in Figure 4.
@@
We further mapped deep features extracted by ResNet-AFFE to observed physical characteristics using UMAP, a nonlinear dimensionality reduction method rooted in graph theory and manifold learning that effectively projects high-dimensional data to low-dimensional spaces for structural analysis. We analyzed 3,833 GRBs detected by Fermi/GBM over the period 2008–2024, concatenating light curves from the triggering detectors of each GRB as model input. Notably, our model uses GAP to aggregate features from the last convolutional layer, enabling compatibility with variable-length data under uniform batch input. This design specifically supports the inclusion of long GRBs with , which would otherwise be excluded by fixed-length constraints. UMAP takes as input the output results from the last convolutional layer of ResNet-AFFE and outputs 2D projections that uncover intrinsic correlations of features between GRBs. Key UMAP hyperparameters including n_neighbors and min_dist were systematically optimized via exhaustive parameter space exploration (over ranges: ; ). Manual validation yielded optimal values of and , achieving a balanced preservation of both global data structure and local neighborhood fidelity—two core advantages of UMAP in high-dimensional signal analysis. The resulting projections shown in Figure 5, visualize the distribution of true positive (TP) and false negative (FN) GRBs, with each GRB’s peak-SNR and encoded directly in the projection to link low-dimensional structure with physical observables. While UMAP effectively reveals the structure of high-dimensional data, its visualization results are sensitive to hyperparameter selection (e.g., n_neighbors and min_dist). To mitigate the subjectivity introduced by manual parameter tuning and ensure the robustness of observed clustering patterns, we implemented a cross-validation procedure.
@@
Following the approach in DownDimGRB_Dimple2023, which targets latent clustering patterns in transient burst events, we further employed the AutoGMM module (autoGMM) to cluster features extracted by the ResNet-AFFE model. AutoGMM is an automated clustering tool that leverages Gaussian Mixture Models (GMMs) to infer the optimal number of clusters, and it assumes data points arise from a mixture of multiple Gaussian distributions, aligning with the statistical properties of GRB feature distributions. The algorithm first performs GMM clustering over a range of candidate cluster counts, then infers the Gaussian distribution parameters and cluster assignments that best fit the data—eliminating subjective manual cluster number selection. The four clusters identified via AutoGMM were visually distinguished using distinct color mappings in the UMAP projections, and their corresponding representative light curves were presented around, as shown in Figure 6. Figure 7 further delineates the spatial distribution of physically distinct GRB subgroups and their proportions within each cluster, including kilonova-associated GRBs (GRB-KN), supernova-associated GRBs (GRB-SN), extended-emission GRBs (GRB-EE), ultra-long GRBs, and short/long GRBs with precursors.

### v2

@@ 1 Introduction
Gamma-ray bursts are among the most energetic phenomena in the universe, releasing enormous amounts of energy in the form of gamma rays over remarkably short time frames. Despite decades of remarkable progress in observations and theoretical investigations, their origins remain elusive (Zhang, 2011; Kumar & Zhang, 2015; Mészáros, 2019; Pe’er, 2024), making GRB identification and classification a critical focus of contemporary astrophysical research (Lü et al., 2010; Zhang et al., 2022; Sun et al., 2023; Wang et al., 2024). The rise of multi-messenger and multi-band astronomy further underscores the need for efficient GRB detection (Margutti & Chornock, 2021; Rudolph et al., 2023): timely and accurate detection enables follow-up observations of afterglows, host galaxy localization, characterization of associated counterparts, and constraints on GRB intrinsic properties.
@@
GRB light curves often exhibit irregular, multi-peaked structures that encode rich physical information about their outbursts. While traditional classification criteria (e.g., duration, hardness) have been widely adopted, emerging evidence highlights the need for additional discriminants to address the ”iceberg effect”—a phenomenon where low-signal-to-noise ratio (low-SNR) GRBs are partially submerged in background noise, disproportionately impacting short-duration events (Lü et al., 2014; Moss et al., 2022). Traditional burst search algorithms identify GRBs by detecting signals in multi-time-bin, multi-energy-band light curves where the SNR exceeds a predefined threshold above the background (Band, 2002; Blackburn et al., 2013; Cai et al., 2021). However, these methods face significant challenges, including accurate background estimation and optimal threshold selection, limiting their effectiveness in identifying faint or complex GRB events. These limitations motivate the development of advanced techniques to enhance GRB detection and classification accuracy.
@@
In recent years, machine learning (ML) has emerged as an experimental tool for GRB detection. For example, Abraham et al. (2021) developed an ML algorithm for automated detection of GRB-like events using AstroSat/CZTI data, combining density-based spatial clustering with dynamic time warping to demonstrate ML’s versatility and robustness in GRB identification. Among ML architectures, convolutional neural networks (CNNs)—a specialized deep learning (DL) framework—excel in pattern recognition tasks (Alzubaidi et al., 2021; Taye, 2023), and have driven a paradigm shift in astrophysical data analysis by autonomously extracting features from multi-dimensional, multi-modal datasets. In GRB research, CNN-based approaches have achieved notable successes. Parmiggiani et al. (2021) developed a real-time CNN pipeline for AGILE-GRID intensity maps, detecting 21 GRBs at 3 significance compared to only two detections via conventional SNR-based searches. Parmiggiani et al. (2023) used a CNN autoencoder to reconstruct background light curves and identify 72 previously uncataloged AGILE GRBs. Additionally, (Parmiggiani et al., 2024) and Crupi et al. (2023) leveraged neural networks to predict background count rates and identify faint, long-duration GRBs.
@@
Notably, GRB identification relies on extracting meaningful features from light curves with complex temporal and spectral characteristics. While traditional 1D-CNNs effectively capture local temporal patterns, their inability to model spectral-temporal correlations limits their utility in GRB analysis—critical GRB features are jointly encoded in both domains (Guidorzi et al., 2016; Tarnopolski & Marchenko, 2021; Zhou et al., 2024). To address this, recent time series classification advances have introduced architectures that explicitly model frequency information, including Wavelet-based decomposition (e.g., T-WaveNet (Liu et al., 2020)), 2D temporal representations (e.g., TimesNet (Wu et al., 2022)), hybrid Fourier-wavelet transformers (e.g., FEDformer (Zhou et al., 2022), WFTNet (Liu et al., 2024)), and adaptive spectral blocks (e.g., TSLANet (Eldele et al., 2024)). These state-of-the-art approaches confirm that spectral-temporal modeling enhances feature extraction, underscoring the need for GRB identification methods that integrate both temporal dynamics and frequency-domain characteristics.
@@
Complementary to classification, dimensionality reduction techniques have provided valuable insights into GRB diversity. For instance, Jespersen et al. (2020) used t-distributed stochastic neighbor embedding (t-SNE) to categorize Swift/BAT GRBs to separate supernova-associated and kilonova-associated GRBs into long and short categories. Garcia-Cifuentes et al. (2023) identified 7 new extended emission (EE) GRB candidates via dimensionality reduction. Zhu et al. (2024) and Negro et al. (2025) used Uniform Manifold Approximation and Projection (UMAP) or t-SNE to reveal GRB clusters that transcend traditional duration-based classification, including distinct subgroups of kilonova-related GRBs. Similarly, Chen et al. (2024) and Chen et al. (2025) demonstrated robust bimodal GRB clustering and refuted intermediate GRB classes using unsupervised ML and dimensionality reduction. Collectively, these studies highlight the utility of such techniques for analyzing burst similarities and refining classification schemes.
@@
Despite these advances, DL models face critical limitations in GRB research. First, data scarcity persists: the largest available GRB dataset (Fermi/GBM, 3900 events) remains insufficient for robust DL training (Zhang et al., 2024), due to limited labeled observations, skewed distribution (underrepresentation of rare, short-duration, and low-SNR GRBs), and narrow coverage of physical properties (e.g., redshift, energy bands, progenitors). While data augmentation has been explored to mitigate this (Zhang et al., 2024; Shorten & Khoshgoftaar, 2019), existing datasets still lack the feature richness needed to avoid overfitting. Second, uncertainty quantification (UQ) is underdeveloped: model uncertainty stems from limited training data, the ”black-box” nature of CNNs, and observational noise (Nemani et al., 2023), yet UQ is critical for distinguishing high-confidence detections from ambiguous candidates—misclassification of which could lead to missed follow-up observations or erroneous inferences about progenitor origins.
@@ 2 Data augmentation and Dataset
The Fermi/GBM trigger system has been operational since 12 July 2008 (first detected GRB: GRB 080714B). For this study, we utilized Fermi/GBM data spanning from 14 July 2008 to 30 June 2024, including 3,905 original GRBs detected by NaI detectors. These events are manually verified and cataloged in the Fermi/GBM Burst Catalog111https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html. Adopting the GRB/non-GRB extraction protocol described in Zhang et al. (2024), we updated the dataset by segmenting samples into 120-second time windows. To ensure complete coverage of the interval (with background included on both sides), 217 GRBs failing this criterion were excluded. For the remaining events, samples were generated based on the number of triggering NaI detectors, resulting in 6,189 primary GRB samples. Non-GRB samples—encompassing modulated background, electronic noise, and potential all-sky source signals—were systematically extracted from detectors’ quiescent (non-trigger) intervals, totaling 108,000 samples. This balance between classes enhances model generalization to non-burst events.
Light curves for all samples were extracted at a 64 ms temporal resolution across 128 energy channels and rebinned into 9 standard energy bands (25–50, 50–100, 50–300, 100–300, 100–500, 100–900, 300–500, 300–900, 500–900 keV). This segmentation aligns with typical GRB photon energy distributions and Fermi/GBM trigger criteria (von Kienlin et al., 2020; Cai et al., 2021), approximating the photon deposition ranges of NaI detectors (without energy response correction). Critically, it preserves energy-dependent features such as burst morphology evolution, light curve shape variations, and spectral lags—enriching the diversity of discriminative features for model training. To optimize input for machine learning, we applied per-band standardization to light curves (scaling features to zero mean and unit variance). This approach preserves spectral information via inter-band flux ratios, empirically outperforming normalization (scaling to [0,1]). While both methods unify feature scales to facilitate deep learning convergence, standardization is preferable for GRB characterization as it retains spectral integrity—essential for distinguishing GRBs from non-GRBs (e.g., via hardness evolution or spectral lags).
@@
Remove 4,648 synthetic samples with peak-SNR below two , as excessive photon subtraction makes these resemble background samples.
@@
Dataset: We partition preprocessed samples into training, validation, and test sets chronologically to ensure temporal generalization. Dataset details are shown in Table 1. This chronological split alleviates temporal overfitting while preserving the data’s inherent flux distribution. We maintain 1:1 positive-to-negative sample ratios across all splits. Crucially, only the training set undergoes data augmentation, while the validation/test sets retain original observations (primary GRB samples and non-GRB samples) for unbiased performance evaluation.
@@ 3.1 Architecture of Neural Networks
Each sample in our dataset consists of light curves spanning 9 energy bands, which are naturally suited to be treated as time series data. This makes a 1D CNN the appropriate choice for feature extraction and classification. Our model adopts ResNet (He et al., 2016) as its backbone, which relies on a sequential stack of four convolutional units (Conv Units) to progressively extract hierarchical features from the input light curves. A key advantage of ResNet is its integration of residual connections, which effectively alleviate the gradient vanishing problem that commonly arises in deep neural networks and enable stable training of deeper architectures. To address the limitation that CNNs lack the ability to distinguish subtle frequency variations, we propose an Adaptive Frequency Feature Enhancement module, AFFE, which is integrated into the last part of each Conv Unit. This module explicitly enhances frequency-domain features to improve the model’s GRB identification capability, and it also adaptively weights and filters frequency components, enabling the model to focus on the most discriminative features for GRB recognition. This design is particularly critical for distinguishing GRBs from background noise, as frequency-domain analysis can uncover patterns that remain obscured in the time domain.
After the convolutional layers (including Conv Units with AFFE modules) complete feature extraction, the extracted high-dimensional features first undergo global average pooling (GAP). This step reduces the number of model parameters while preserving the representative information of features across the temporal dimension, thereby mitigating the risk of overfitting. A dropout layer with a 50% dropout rate is then inserted between the GAP layer and the subsequent fully connected (FC) layers. During training, this layer randomly deactivates half of the neurons to prevent the model from over-relying on specific features, further suppressing overfitting. Following the dropout layer, two FC layers map the pooled features to the final classification space and output the probability of the sample belonging to the GRB or non-GRB category. The overall architecture of our ResNet-based model integrated with AFFE modules is shown in Figure 3, which includes the order of layers and the data flow.
@@ 3.2 Model training and analysis
Model parameters are initialized using the truncated normal distribution proposed by He et al. (2015). A detailed analysis of hyperparameter choices-including batch size, learning rate, and number of convolutional layers is conducted to evaluate their impact on the test set, following the approach in Ma et al. (2023). Additional hyperparameters, including the number of convolutional kernels per unit, the size of the convolutional kernels, and the neuron counts in the fully connected (FC) layers, were also optimized. The training configuration is as follows: batch size is set to 1024; cross-entropy loss is used to measure the discrepancy between predicted and true labels; and the Adam optimizer Kingma & Ba (2014) is adopted to dynamically update model parameters and minimize loss during training. The initial learning rate is , with a scheduling mechanism that reduces the rate by half if validation loss does not decrease for 10 consecutive epochs. To prevent overfitting, early stopping terminates training if validation accuracy plateaus for 20 consecutive epochs. The patience value was selected after testing ranges of [10, 20, 40] epochs, and the model parameters from the epoch with the highest validation accuracy are retained as the final optimal parameters. The result of the optimizing choices are highlighted in bold in Table 2. The efficacy of selected hyperparameters is further validated by training curves, which demonstrate stable convergence and appropriate model capacity for the classification task.
@@
We conducted a comparative analysis involving the baseline ResNet architecture that is commonly used for time series data, and ResNet variants integrated with existing feature enhancement modules that including the SE (Squeeze-and-Excitation block; Hu et al. (2018)), CBAM (Convolutional Block Attention Module; Woo et al. (2018)), ECA (Efficient Channel Attention module; Wang et al. (2020)), and ASB (Adaptive Spectral Block; Eldele et al. (2024) modules), and our proposed ResNet-AFFE model. All models were trained and tested on our dataset using hyperparameters consistent with those reported in the original literature for each comparative module. This ensures a fair performance comparison, and the results of this comparison (for ResNet variants and our ResNet-AFFE model) on the test set are presented below.
The models were implemented using the PyTorch framework, with training conducted on a single NVIDIA RTX-4090 GPU. Performance was evaluated using four standard classification metrics: Accuracy, Precision, Recall, and F1-score, consistent with the evaluation methodology reported in Zhang et al. (2024). Accurately quantifying these metrics and their associated prediction uncertainties is critical, as they quantify confidence in model predictions and provide a scientific basis for subsequent decision-making. During testing, we achieved robust uncertainty quantification for deep learning predictions via Monte Carlo Dropout (MCD, for references see Gal & Ghahramani (2015); Basora et al. (2025)). In this procedure, all model parameters are kept fixed, and dropout layers between FC layers remain activated. For each forward pass, the dropout rate is randomly sampled from a uniform distribution over the interval [0.1, 0.5]. This stochastic forward pass is repeated 1000 times, which allows computation of the mean and standard deviation for each metric—thereby effectively estimating predictive uncertainty. The performance and uncertainty of all models are summarized in Table 3.
@@
Enhanced time domain feature extraction module: a Hu et al. (2018), b Woo et al. (2018), c Wang et al. (2020).
Enhanced frequency domain feature extraction module: d Eldele et al. (2024).
@@
To examine whether ResNet-AFFE prioritizes burst-specific temporal signatures (e.g., intervals or energy-dependent flux peaks) over noise, we used Gradient-weighted Class Activation Mapping (Grad-CAM)—a mainstream visualization technique for addressing deep learning’s ”black box” problem. Grad-CAM generates gradient-weighted heatmaps to identify features most influential to model predictions (Selvaraju et al., 2017), enhancing interpretability. For time-series light curves, these heatmaps highlight temporal segments driving classification outputs, replacing ambiguous ”feature importance” with intuitive visual evidence. We employ Grad-CAM to decode our model’s decision-making process for GRB classification. Comparative visualizations of discriminative temporal features between the baseline ResNet and ResNet-AFFE are presented in Figure 4.
@@
We further mapped deep features extracted by ResNet-AFFE to observed physical characteristics using UMAP, a nonlinear dimensionality reduction method rooted in graph theory and manifold learning that effectively projects high-dimensional data to low-dimensional spaces for structural analysis. We analyzed 3,833 GRBs detected by Fermi/GBM over the period 2008–2024, concatenating light curves from the triggering detectors of each GRB as model input. Notably, our model uses GAP to aggregate features from the last convolutional layer, enabling compatibility with variable-length data under uniform batch input. This design specifically supports the inclusion of long GRBs with , which would otherwise be excluded by fixed-length constraints. UMAP takes as input the output results from the last convolutional layer of ResNet-AFFE and outputs 2D projections that uncover intrinsic correlations of features between GRBs. Key UMAP hyperparameters including n_neighbors and min_dist were systematically optimized via exhaustive parameter space exploration (over ranges: ; ). Manual validation yielded optimal values of and , achieving a balanced preservation of both global data structure and local neighborhood fidelity—two core advantages of UMAP in high-dimensional signal analysis. The resulting projections, shown in Figure 5, visualize the distribution of true positive (TP) and false negative (FN) GRBs, with each GRB’s peak-SNR and encoded directly in the projection to link low-dimensional structure with physical observables. While UMAP effectively reveals the structure of high-dimensional data, its visualization results are sensitive to hyperparameter selection (e.g., n_neighbors and min_dist). To mitigate the subjectivity introduced by manual parameter tuning and ensure the robustness of observed clustering patterns, we implemented a cross-validation procedure.
@@
Following the approach in Dimple et al. (2023), which targets latent clustering patterns in transient burst events, we further employed the AutoGMM module (Athey et al., 2019) to cluster features extracted by the ResNet-AFFE model. AutoGMM is an automated clustering tool that leverages Gaussian Mixture Models (GMMs) to infer the optimal number of clusters, and it assumes data points arise from a mixture of multiple Gaussian distributions, aligning with the statistical properties of GRB feature distributions. The algorithm first performs GMM clustering over a range of candidate cluster counts, then infers the Gaussian distribution parameters and cluster assignments that best fit the data—eliminating subjective manual cluster number selection. The four clusters identified via AutoGMM were visually distinguished using distinct color mappings in the UMAP projections, and their corresponding representative light curves were presented around, as shown in Figure 6. Figure 7 further delineates the spatial distribution of physically distinct GRB subgroups and their proportions within each cluster, including kilonova-associated GRBs (GRB-KN), supernova-associated GRBs (GRB-SN), extended-emission GRBs (GRB-EE), ultra-long GRBs, and short/long GRBs with precursors.